# Relaxation time approximation revisited and pole/cut structure in retarded correlators

### v2

@@ V The novel relaxation time approximation
which is referred to as the novel RTA in [27], and is also known as the mutilated operator in [34, 36, 37] (see chapter V of [19] for an earlier discussion). Here $|\psi_{n}\rangle$ denotes the orthonormal eigenfunctions of $\mathcal{L}_{0}$ with zero eigenvalues, and $\gamma_{6}$ represents the smallest positive eigenvalue with the dimension of $[E]$. Thus we recover the collision invariance: $\mathcal{L}_{0}|\psi_{n}\rangle=0,n=1,\cdots 5$. To relate it to the traditional RTA, one needs to identify the relaxation time as $\tau_{R}\equiv\frac{1}{\gamma_{6}}$. When $\mathcal{L}_{0}$ acts on other eigenfunctions, it results in $\mathcal{L}_{0}|\psi_{n}\rangle=\gamma_{6}|\psi_{n}\rangle$ for $n>5$, collapsing all positive eigenvalues into the smallest positive one. This is why the model in Eq.(36) is sometimes referred to as ”mutilated” [19].
Macroscopic conservation laws are inherently restored when microscopic conservation laws, specifically the collision invariance of the collision operator, are respected in the model’s construction, thereby simultaneously fixing the basic flaws. The novel RTA provides the flexibility to adjust the matching conditions used in kinetic theory, a feature not present in the traditional RTA as discussed in [27, 34]. This flexibility is particularly advantageous when discussing hydrodynamic frame dependence, such as in the first-order causal theory of the BDNK type [38, 39, 40, 41].
@@
$-\mathcal{L}_{0}\chi(p)=-\frac{1}{\tau_{R}}\left[\chi(p)-\frac{\left(1,\chi(p)\right)}{\left(1,1\right)}-P^{(0)}_{1}\frac{\left(1,P_{1}^{\left(0\right)}\chi(p)\right)}{\left(1,P_{1}^{\left(0\right)}P_{1}^{\left(0\right)}\right)}-p^{\left\langle\mu\right\rangle}\frac{\left(1,p_{\left\langle\mu\right\rangle}\chi(p)\right)}{(1/3)\left(1,p_{\left\langle\nu\right\rangle}p^{\left\langle\nu\right\rangle}\right)}\right].$ (37)
@@ VI Summary and outlook
In this paper, we revisit the widely-used relaxation time approximation within the linearized Boltzmann equation. According to the mathematical analysis on the eigenspectrum of linearized Boltzmann collision operator $\mathcal{L}_{0}$, the RTA model is justified only for hard interactions, thereby ruling out the energy-dependent parametrization in Eq.(8). The consideration is grounded on mathematical aspects, and Eq.(8) can effectively serve as a convenient parameterized model. Furthermore, Eq.(8) sheds light on the redefinition of linearized collision operator denoted as $\mathcal{L}_{\alpha}$. We demonstrate that, using only physical arguments, one can related the eigenspectrum properties of the redefined $\mathcal{L}_{\alpha}$ to the softness of interactions through the available mathematical analysis on $\mathcal{L}_{0}$. This, in our view, is nontrivial, as intuitive physical considerations can yield meaningful mathematical insights without the need for sophisticated analysis. Thus far, we have completed the summary of Q1. We also provide a derivation of the novel RTA to restores collision invariance by truncating the linearized collision operator, rather than by adding counter terms.
@@
Regarding Q2, we find that focusing on $\mathcal{L}_{0}$ and its eigenspectrum properties is the simplest approach. When interactions are hard, the RTA is well-defined, and the analytical properties of retarded correlators, as detailed in [7], are applicable. There is a gap between the branch-cut lines and the real axis $\text{Im}\,\omega=0$. Therefore, the gapless hydrodynamic modes are well-defined low energy degrees of freedom when $k$ is small. However, according to mathematical derivations, relativistic interactions are often soft, leading us to focus on the alternative scenario in most cases. In scenarios with soft interactions, the RTA is no longer well-justified, and the dominating long-lived non-analytical structure turns into the branch-cut or the non-hydrodynamic modes. Our conclusion is consistent with the previous related studies with the comparison details elaborated in the main text. Note that if particles constituting the system are massive or the perturbations are inhomogeneous, the non-analytical structures would be more complicated and richer.
@@
There are possible extensions to the present research. As mentioned in III.3, we can establish the dictionary between various field theories and the eigenspectrum structure within their linearized kinetic description, which can help us to determine the dominating non-analytical structure and get all possible RTA models. The former concerns with the properties of the retarded correlation function of many-body systems, and the latter offers a solid theoretical basis for RTA if any. We believe that the application of RTA to a justified system is theoretically more consistent than that without justification, which is one of the motivations for this script. Another relevant extension is massive transport within the RTA. One may think of it as a trivial follow-up but nonzero mass do contribute nontrivially to our discussions as seen in Eq.(32) and the second condition mentioned in III.3. There is one more hint for its nontriviality: if one recalculate the retarded correlators for massive particles, the denominator turns complicated due to the nonzero mass and the whole integral can not be worked out analytically in contrast to massless case, indicating that there might be distinct non-analytical structures compared to Romatschke’s work, as discussed in Sec.IV. Also, the numerical calculation for massive scalar $\phi^{4}$ theory with leading-order interaction is also in progress. Last but not least, almost all related researches on the properties of retarded correlators base themselves on the linearized kinetic description. It would be interesting to seek the impacts of the nonlinear structure contained in the complete kinetic description, e.g. Boltzmann equation, on the present conclusions.

### v3

@@ V The novel relaxation time approximation
which is referred to as the novel RTA in [21, 32], and is also known as the mutilated operator in [40, 41, 42] (see chapter V of [22] for an earlier discussion). Here $|\psi_{n}\rangle$ denotes the orthonormal eigenfunctions of $\mathcal{L}_{0}$ with zero eigenvalues, and $\gamma_{6}$ represents the smallest positive eigenvalue with the dimension of $[E]$. Thus we recover the collision invariance: $\mathcal{L}_{0}|\psi_{n}\rangle=0,n=1,\cdots 5$. To relate it to the traditional RTA, one needs to identify the relaxation time as $\tau_{R}\equiv\frac{1}{\gamma_{6}}$. When $\mathcal{L}_{0}$ acts on other eigenfunctions, it results in $\mathcal{L}_{0}|\psi_{n}\rangle=\gamma_{6}|\psi_{n}\rangle$ for $n>5$, collapsing all positive eigenvalues into the smallest positive one. This is why the model in Eq.(34) is sometimes referred to as ”mutilated” [22].
Macroscopic conservation laws are inherently restored when microscopic conservation laws, specifically the collision invariance of the collision operator, are respected in the model’s construction, thereby simultaneously fixing the basic flaws. The novel RTA provides the flexibility to adjust the matching conditions used in kinetic theory, a feature not present in the traditional RTA as discussed in [43, 32, 44, 40, 45]. This flexibility is particularly advantageous when discussing hydrodynamic frame dependence, such as in the first-order causal theory of the BDNK type [46, 47, 48, 49].
@@
$\mathcal{L}_{0}\chi(p)=-\frac{1}{\tau_{R}}\left[\chi(p)-\frac{\left(1,\chi(p)\right)}{\left(1,1\right)}-P^{(0)}_{1}\frac{\left(1,P_{1}^{\left(0\right)}\chi(p)\right)}{\left(1,P_{1}^{\left(0\right)}P_{1}^{\left(0\right)}\right)}-p^{\left\langle\mu\right\rangle}\frac{\left(1,p_{\left\langle\mu\right\rangle}\chi(p)\right)}{(1/3)\left(1,p_{\left\langle\nu\right\rangle}p^{\left\langle\nu\right\rangle}\right)}\right].$ (35)
@@ VI Summary and outlook
In this paper, we revisit the widely used relaxation time approximation within the linearized Boltzmann equation. According to the mathematical analysis on the eigenspectrum of linearized Boltzmann collision operator $\mathcal{L}_{0}$, the RTA model is justified only for hard interactions, thereby ruling out the energy-dependent parametrization in Eq.(8). The consideration is grounded on mathematical aspects, and Eq.(8) can effectively serve as a convenient parameterized model. Furthermore, Eq.(8) sheds light on the redefinition of linearized collision operator denoted as $\mathcal{L}_{\alpha}$. We also provide a derivation of the novel RTA to restore the collision invariance by truncating the linearized collision operator, instead of by adding counter terms.
@@
For discussing the non-analytical structures within the retarded correlators, we find that focusing on $\mathcal{L}_{0}$ and its eigenspectrum properties is the simplest approach. When interactions are hard, the RTA is well-defined, and the analytical properties of retarded correlators, as detailed in [7], are applicable. There is a gap between the branch-cut lines and the real axis $\text{Im}\,\omega=0$. Therefore, the gapless hydrodynamic modes are well-defined low-energy degrees of freedom when $k$ is small. However, according to mathematical derivations, relativistic interactions are often soft, leading us to focus on the alternative scenario in most cases. In scenarios with soft interactions, the RTA is no longer well-justified, and the dominating long-lived non-analytical structure turns into the branch-cut or the non-hydrodynamic modes. Our conclusion is consistent with the previous related studies with the comparison details elaborated in the main text. Note that if the particles constituting the system are massive or the perturbations are inhomogeneous, the non-analytical structures would be more complicated and richer.
@@
There are possible extensions to the present research. As mentioned in III.4, we can establish a dictionary between various field theories and the eigenspectrum structure within their linearized kinetic description. This can help us determine the dominant non-analytical structure and derive all possible RTA models. The former concerns the properties of the retarded correlation function of many-body systems, while the latter provides a solid theoretical basis for RTA, if any. We believe that applying RTA to a justified system is theoretically more consistent than doing so without justification, which is one of the motivations for this work. Additionally, almost all related research on the properties of retarded correlators is based on the linearized kinetic description. It would be interesting to explore the impacts of the nonlinear structure contained in the complete kinetic description, such as the Boltzmann equation, on the present conclusions.
Acknowledgments
This work is supported by NSFC under grant No.12505149.